OpenStudy (anonymous):

Given the formula E = IR what is the formula for R? A. R = E ÷ I B. R = EI C. R = I ÷ E D. R = IE

OpenStudy (anonymous):

E = IR divide by I what do you get

OpenStudy (jiteshmeghwal9):

divide i both sides :)

OpenStudy (anonymous):

i still dont understnad this is one thing i really dont get

OpenStudy (anonymous):

a*b=c*d; solve for c a*b/d=c*d/d ab/d = c

OpenStudy (anonymous):

never mind i got r=e/i

OpenStudy (anonymous):

yes

OpenStudy (jiteshmeghwal9):

we said the divide i both sides\[\Huge{\frac{E}{I}={IR \over I}}\]

OpenStudy (anonymous):

oh okay thank you all

OpenStudy (jiteshmeghwal9):

yw :)